OpenStudy (anonymous):

For Josh's science project he is filling a mountain with fake lava liquid. The mountain he created was in the shape of a square pyramid with a base length of 12 inches. If 432 square inches of liquid fits into the pyramid, what is the pyramid's height?

OpenStudy (mathstudent55):

For a pyramid, the volume is \( V = \dfrac{1}{3}Bh \) where B = area of the base. For a square, the area is \(A = s^2 \) where s = length of the side of the square.

OpenStudy (mathstudent55):

The amount 432 is in cubic inches or square inches?

OpenStudy (anonymous):

square inches

OpenStudy (anonymous):

@mathstudent55 would the hieght be 36

OpenStudy (mathstudent55):

How can you fill something with an amount in square inches? Square inches is a unit of area. If you're filling a pyramid, you need an amount in cubic inches, a unit of volume.

OpenStudy (mathstudent55):

Is the pyramid being filled with fake lava or covered with fake lava?

OpenStudy (anonymous):

filled

OpenStudy (mathstudent55):

Is 432 supposed to be in cubic inches and is the volume of the pyramid, or is 432 supposed to be in square inches an is the surface area of the pyramid?

OpenStudy (mathstudent55):

If the pyramid will be filled, then it has to be 432 cubic inches, not 432 square inches, of fake lava.

OpenStudy (mathstudent55):

\(V = Bh \) \( 432 ~in.^3 = 12^2 ~in.^2 h\) \(h = \dfrac{432~in.^3}{144 ~in.^2} = h \) \(h = 3 ~in. \)

OpenStudy (anonymous):

0hk thanks for explainig cause i would have put sq inches

OpenStudy (mathstudent55):

wlcm